OpenStudy (anonymous):

What two consecutive positive integers whose product is 110?

OpenStudy (anonymous):

what's 10 x 11? First look at all the numbers that are multiplied to get 110. Then see which ones are consective to each other.

OpenStudy (anonymous):

how do you know which are consective?

OpenStudy (anonymous):

(x)(x + 1) = 110 x^2 + x - 110 = 0 (x - 10)(x + 11) = 0 x = 10, -11 positive so x = 10 10 and 11

OpenStudy (anonymous):

consecutive means going in order like 1, 2, 3, 4, 5, etc